OpenStudy (anonymous):

In 1991, the life expectancy of males in a certian country was 66.1 years. In 1997, it was 69.7 years. Let E represent the life expectancy in year T and let T represent the number of years since 1991. Linear function E(t)=?t+? predict life expectancy of males in 2005

OpenStudy (anonymous):

E(t) = 0.6 t + 66.1

OpenStudy (dumbcow):

slope = (69.7-66.1)/6 = 0.6 E(t) = 0.6t + 66.1 E(14) = 8.4 +66.1 = 74.5